The length of a new rectangular field is 7 yards longer than triple the width. If the perimeter of the rectangular playing field is 470 yards what are its dimensions?
Answer by checkley77(12844)

L=3W+7
P=2L+2W
470=2(3W+7)+2W
470=6W+14+2W
470=8W+14
8X=470-14
8W=456
W=456/8
W=57 YDS. IS THE WIDTH.
L=3*57+7=178 YDS. FOR THE LENGTH.
PROOF:
470=2*178+2*57
470=356+114
470=470